OpenStudy (babyd):

and 3 and 1/5 again converting from fraction to percent

jimthompson5910 (jim_thompson5910):

\[3\frac{1}{5}=3+\frac{1}{5}=3+0.2=3.2\] Now multiply by 100 to get 320%